OpenStudy (blondey_315):

solve for x: |x+9|=17 solve for n: |3n+5|=7

OpenStudy (anonymous):

x = 8 17- 9 = 8 x = 8 n = 2/3

OpenStudy (anonymous):

When solving for absolute value take the equation and make it equal the value and the negative value of the solution...then solve for x.

OpenStudy (anonymous):

Thus the first equation we need to solve x+9 = 17 and x+9 = -17

myininaya (myininaya):

Let's no just give all the solutions to these. @blondey_315 needs help not just answers.

myininaya (myininaya):

We should let her try to answer some of them.

OpenStudy (anonymous):

I just finished a lesson on this

OpenStudy (anonymous):

for |3n+5|=7 |3n+5|=7 (3n+5)^2 = 7^2 3n+5 = + - sqrt(7^2) 3n + 5 = + - 7 So 3n + 5 = 7 n = (7-5)/3 = 2/3 or 3n+5 = -7 3n = -7-5 n =-12/3 n = -4

OpenStudy (blondey_315):

Yes please! Those answers werent on the answer choices either

OpenStudy (anonymous):

Thus the first equation we need to solve x+9 = 17 and x+9 = -17

OpenStudy (anonymous):

What are the answer choices? That would help us out a bit.

OpenStudy (anonymous):

When solving for absolute value take the equation and make it equal the value and the negative value of the solution...then solve for x.

OpenStudy (blondey_315):

Thank you!

OpenStudy (anonymous):

sorry for |x+9|=17 (x+9)^2 = 17^2 x+9 = + - sqrt(17^2) x+9 = + - 17 so x = 17 - 9 = 8 or x = -17-9 = -26

OpenStudy (anonymous):

x = -16 or x = -2 for first absolute value equation.

OpenStudy (anonymous):

SO the answer is X = 8 or X = -26 and n = 2/3 or n = -4

OpenStudy (anonymous):

Why are you making this more difficult than it has to be?

OpenStudy (anonymous):

You are welcome

OpenStudy (anonymous):

Who?

OpenStudy (blondey_315):

@pip_pip123 the answer choices for the first one is (8, 26) (8, -26) (-8, 26) (-8, -26)

OpenStudy (anonymous):

Oh this is much easier

OpenStudy (anonymous):

(8,-26)

OpenStudy (anonymous):

|a+b| = c thus a+b=c and a+b = -c

OpenStudy (anonymous):

@blondey_315 : get it baby?